molarity of

3.0 g of KOH in 3.5 L of KOH solution
and
6.05 g of NaCl in 500. mL of NaCl solution
1
Expert's answer
2017-12-11T07:36:41-0500
Solution
1) Find the amount of KOH in the solution:
ν = = 0.054 (mol)
Find the molarity:
C = = 0.015 (mol/L).
2) Find the amount of NaCl in the solution:
ν = = 0.103 (mol)
Find the molarity:
C = = 0.206 (mol/L).
Answer
Molarity of KOH solution: 0.015 (mol/L).
Molarity of NaCl solution: 0.206 (mol/L).